Bituminous coal is a type of coal that is composed of carbon with small amounts of hydrogen, oxygen, nitrogen, and sulfur. The formula for a sample of bituminous coal is 

C137H97O9NS. Calculate the amount of carbon in 65.0 kg of the sample.


1
Expert's answer
2022-03-17T03:34:03-0400

The molecular mass of C137H97O9NS is 1933.30

The percentage of carbon in it is ("137*12.01\/1933.30" )*100% = 85.11% (or 0.8511)

65.0 * 0.8511 = 55.32 (kg)